OpenStudy (anonymous):

What is 7.8 × 105 m/h in kilometers per day? A. 3.24 × 1010 km/day B. 3.24 × 108 km/day C. 1.872 × 106 km/day D. 1.872 × 104 km/day

OpenStudy (anonymous):

please explane how you got the answer too

OpenStudy (anonymous):

will become fan

OpenStudy (anonymous):

Is that meters/hour into km/day?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

are you still there

OpenStudy (anonymous):

How many hours in a day? How many meters in a kilometer?

OpenStudy (anonymous):

24 1000 i think b am i corect

OpenStudy (anonymous):

are you still there

OpenStudy (anonymous):

\[\frac{7.8\times10^5\text{ m}}{1\text{ hr}}\times\frac{24\text{ hr}}{1\text{ day}}\times\frac{1\text{ km}}{1000\text{ m}}\] In terms of scientific notation: \[\frac{7.8\times10^5\text{ m}}{1\text{ hr}}\times\frac{2.4\times10^1\text{ hr}}{1\text{ day}}\times\frac{1\text{ km}}{1\times10^3\text{ m}}\] The units will cancel out nicely: \[\frac{7.8\times10^5}{1}\times\frac{2.4\times10^1}{1\text{ day}}\times\frac{1\text{ km}}{10^3}\] Reduce the powers of 10: \[(7.8\times2.4)\times10^3\frac{\text{km}}{\text{day}}\]

OpenStudy (anonymous):

And you have \[7.8\times2.4=18.72=1.872\times10^1\] so you have \[(1.872\times10^1)\times10^3\frac{\text{km}}{\text{day}}=1.872\times10^4\frac{\text{km}}{\text{day}}\]

OpenStudy (anonymous):

Make sense?

OpenStudy (anonymous):

your answer is not an option though

OpenStudy (anonymous):

Have you considered checking D?